OpenStudy (anonymous):

In a town, there are 170 teenagers that own portable media players. This is at least 2/3 of all the teenagers in the town. What is the maximum number of teenagers who live in the town? Please explain. Don't just give me the answer. Thank you.

OpenStudy (anonymous):

Crying, i have to retype it, one sec

OpenStudy (anonymous):

Oh jeeze!! I'm sorry lol. But thank you for the help. It means a lot.

OpenStudy (anonymous):

Okay, so notice that what you're asked is to find the total population, so lets denote that as our variable "x", alrighty, now lets look at what we know, we are told that 2/3 of our population "x" own portable media players, and that 2/3 of this population is equivalent to 170, writing this out mathematically, we have that\[170=\frac{ 2 }{ 3 }x\]since 2/3rds of the population is given to us as the value 170 Note: since it says the MAXIMUM, that means that if you get (and you will) a decimal answer, you want to round it up, because you can't have a fraction of a teenager (i sure hope so)

OpenStudy (anonymous):

to find your answer you would use this: 170 = 2/3(x) x is the total population. 170 is at least 2/3 of the total population. round your answer up since this is regarding number of people.

OpenStudy (anonymous):

So I multiply 170 by .67?

OpenStudy (anonymous):

No, you should divide 170 by .67 because 2/3(x) = 170 or .67(x) = 170 so 170/(2/3) = x or 170/.67 = x

OpenStudy (anonymous):

Ohhh ok!! So my answer would be x = 254 teenagers since I got a decimal of 253.7?

OpenStudy (anonymous):

Yes! Very good :)

OpenStudy (anonymous):

Sweet!! Thank you so much!!

OpenStudy (anonymous):

You're welcome :)